OpenStudy (anonymous):

MEDAL Which value of x makes the following equation true? * 2(x + 8) - 2 = -5x

OpenStudy (jango_in_dtown):

2(x + 8) - 2 = -5x implies 2x+16-2=-5x implies 2x+5x=-14 implies 7x=-14 implies x=-2

OpenStudy (ddos_dragon):

Beat me to it :P

OpenStudy (ddos_dragon):

Oh look, a Sabu impersonator! :3